# Neuromorphic Computing with Microfluidic Memristors

### v1

@@ IV Coupling Memriki oscillators
Figure 4: Circuit diagram of a XOR-pre-gate (a) and a NAND-pre-gate (c) constructed by combining three Memriki oscillators of Fig. 2 (here represented by the three boxes), either only coupled capacitively (a) or with two additional operational amplifiers (c). Time traces of the two input voltages (blue and orange) representing the four consecutive binary states 10, 00, 01, 11. The resulting quasi-binary output voltage (green), which is either quiescent or oscillating, is shown in (b) for circuit (a), revealing an XOR behavior, and in (d) for circuit (c), exhibiting a NAND behavior. The corresponding truth tables are shown alongside the circuit diagrams.
To use the output of one logic gate as the input for another, the output must be compatible with the input requirements, in this case static signals of $0\,\mathrm{V}$ and $1\,\mathrm{V}$. To achieve a proper NAND gate, we must transform the oscillatory output of our NAND-pre-gate into a constant output of $1\,\mathrm{V}$ while its quiescent (0) output for the 11 input remains unaffected. This transformation is achieved by passing the output voltage of our NAND-pre-gate through the coupler shown in the circuit diagram of Fig. 5. In this coupler, the output voltage first passes through operational amplifier OA1, which strengthens the signal and prevents drawing any current from the input gates. The amplified signal is then passed through a full-wave bridge rectifier composed of four diodes, followed by time-averaging with a low-pass filter with time constant $R_{LP}C_{LP}=0.5\,\mathrm{ms}$. Finally, the signal is amplified by AO2 to achieve the desired $1\,\mathrm{V}$ logic level.
Figure 5: Coupler for converting oscillations into constant logic levels. The oscillatory input voltage difference between $U_{\mathrm{in}}^{+}$ and $U_{\mathrm{in}}^{-}$ is first amplified by the left operational amplifier, then rectified by a full-wave bridge rectifier composed of four diodes. The rectified signal is subsequently smoothed by a low-pass filter consisting of $R_{\mathrm{LP}}$ and $C_{\mathrm{LP}}$ and finally amplified to the desired logic level by the right operational amplifier. The operational amplifiers in this circuit are modeled by Eq. 3, with a maximum voltage of $U_{\mathrm{cc}}=1\,\mathrm{V}$.
@@
where $G$ is the gain factor (for small inputs), set to $G=15$ for $\mathrm{OA}_{1}$ and to $G=35$ for $\mathrm{OA}_{2}$. The parameter $\Delta U=1.0\,\mathrm{V}$ defines the range of the amplifier, allowing amplification up to ±$\pm 1\,\mathrm{V}$. The low-pass RC filter consists of a resistance $R_{\mathrm{LP}}=10^{-2}\,\mathrm{g_{0}}^{-1}$ and a capacitance $C_{\mathrm{LP}}=5\times 10^{-2}\,\mathrm{g_{0}s}$, while the diodes are modeled with a saturation current of $0.3\,\mathrm{g_{0}V}$ and an emission factor $\eta=1$ [42].
With the addition of this coupler circuit, the NAND-pre-gate becomes a true NAND-gate.
@@ V Combining gates
The NAND gate is well known to be functionally complete, which means that any other logic gate can be constructed from it. In particular, an OR gate can be constructed using three NAND-gates, as shown in Fig. 6(a). The Boolean logic for the OR gate follows $\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{1}}\cdot\overline{U_{\mathrm{in}}^{2}\cdot U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}}\cdot\overline{U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}}}+\overline{\overline{U_{\mathrm{in}}^{2}}}=U_{\mathrm{in}}^{1}+U_{\mathrm{in}}^{2}$, where $\cdot$ denotes the logic AND, $+$ the logic OR, and an overline indicates negation. In our circuits, each input signal $U_{\mathrm{in}}^{1}$ and $U_{\mathrm{in}}^{2}$ is passed through the combined gates twice to ensure independent noise sources for all our Memriki gates. If identical signals were used, the differential input to the final gate would cancel out, preventing the oscillations from self-exciting. To further aid self-excitation, we also add Gaussian noise with a standard deviation of $0.003\,\mathrm{V}$ to the inputs of the last gate.
@@
As is well known, the circuit for the OR-gate can be reused to realize an AND-gate; only the inputs need to be rearranged such that both input gates now receive $U_{\mathrm{in}}^{1}$ and $U_{\mathrm{in}}^{2}$, as shown in Fig. 6(b). The logic for the AND gate is as follows: $\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}\cdot\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}+\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}=U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}$. The resulting input-output relation for this AND-gate is shown in Fig. 6(d), confirming that it functions correctly as an AND gate. The dynamic response to slowly varying inputs is comparable to that of the OR gate.
Figure 6: Three combined NAND gates (see text) form (a) an OR gate and (b) an AND-gate. Time traces of the input voltages (blue and orange), representing four consecutive binary states, and the corresponding output voltage (green) for (c) the OR gate and (d) the AND-gate, along with the corresponding truth tables.
To test the repeatability and reproducibility of the gates we need to verify whether they perform as expected over many cycles of the (weakly noisy) input and output sequences. For our continuous system, with output voltages spanning the full range between zero and one volt when the input signal switches, we need to define boundaries separating high- and low-level signals during specific time intervals. Here we set the time interval as the duration between the switching of the input signals and consider an output larger than $0.66\,\mathrm{V}$ as a digital 1 and below $0.33\,\mathrm{V}$ a digital 0. To evaluate the pre-gates, we connect them to the coupling circuit as described above to ensure they function as proper gates. Using these definitions, we tested 100 cycles of all possible input combinations and obtained the following accuracies. Individual XOR-gates operate with accuracy of $\geq 99\%$, NAND-gates with $\geq 99\%$, the combined OR-gate $\geq 99\%$, and the combined AND-gate with $\geq 95\%$ of the time.
@@ VI Conclusion
We designed microfluidic iontronic circuits, mainly based on the combination of volatile memristors and Shinriki-like oscillators, that can perform the non-linear XOR classification task and act as the functional complete NAND logic gate. We showed that these NAND-gates are robust and can be combined to form other gates, especially OR and AND. A key feature of these circuits is their exploitation of noise, here set at the level of $3\,\mathrm{mV}$ on the scale of $1\,\mathrm{V}$ for the signals. This noise level allows for self-excitation of the oscillators from a quiescent (0) to an oscillating (1) state, a feature that stems from the intrinsic non-linearity of the “Memriki” oscillators from which our circuits were composed. We hope that our findings will inspire future explorations of experimental realizations of memristive iontronic devices. While we have not investigated the energy efficiency of our circuits and logic gates in any detail yet, we do not expect their current state to be more energy efficient than present-day silicon technology. In this sense, this work serves as a proof of concept for a potentially versatile iontronic platform for computing, based on the same aqueous electrolyte as mammalian brains.
@@ VII Methods
We stress that this explicit expression for $g_{\infty}(U)$ stems from an (approximate) analytic solution to the fully microscopic Poisson-Nernst-Planck-Stokes equations in the long-channel limit. The dependence of the conductance on $U$ is a direct consequence of electro-osmotic flow, characterized by the dimensionless parameter $\mathrm{Pe}(U)$. This effect relies on the presence of a nonzero surface charge $\sigma$ (and hence a nonzero zeta potential $\psi$ on the channel walls), as becomes explicit by the analytic relation $\mathrm{Pe}(U)=\frac{-eR_{b}}{k_{B}TR_{t}w}U$, where $w=\frac{eD\eta}{k_{B}T\varepsilon\psi}\approx-9.6$ represents the ratio of ionic mobility $D/k_{B}T$ to osmotic mobility $\varepsilon\psi/(e\eta)$. Here $D=1.75\,\mathrm{nm}^{2}/\mathrm{ms}$ is the diffusion coefficient of the monovalent cations and anions in water, $k_{B}$ denotes the Boltzmann constant, $T=293.15\,\mathrm{K}$ is the absolute temperature, and $\varepsilon=80.2\epsilon_{0}\approx 7.1\times 10^{-8}\,\mathrm{F/m}$ is the dielectric constant, and $\eta=1.01\,\mathrm{mPas}$ is the shear viscosity of water. We define the steady zero-field conductance as $g_{0}=\frac{\pi R_{t}R_{b}}{L}\frac{2\rho_{b}e^{2}D}{k_{B}T}$, where $\rho_{b}$ is the ionic bulk concentration, set to $10\,\mathrm{mM}$. The asymmetry parameter is given by $\Delta g=\frac{-2w\Delta R}{R_{b}\mathrm{Du}}\approx-3.59$, where the Dukhin number is $\mathrm{Du}=\sigma/(2\rho_{b}R_{t})\approx-0.25$, with $\sigma=-0.0015\,e/\mathrm{nm}^{2}$ the surface charge density on the channel walls. The corresponding zeta potential is $\psi=\frac{2k_{B}T}{e}\mathrm{asinh}(2\pi\lambda_{D}\lambda_{B}\sigma/e)\approx-10\,\mathrm{mV}$, where $\lambda_{D}=\sqrt{\varepsilon k_{B}T/(2e^{2}\rho_{b})}\approx 30\,\mathrm{nm}$ is the Debye length and $\lambda_{B}=e^{2}/(4\pi\varepsilon k_{B}T)\approx 0.71\,\mathrm{nm}$ is the Bjerrum length.

### v2

@@ IV Coupling Memriki oscillators
Figure 5: Circuit diagrams of a XOR-pre-gate (a) and a NAND-pre-gate (c) constructed by combining three Memriki oscillators from Fig. 2 (represented here by the three boxes), either coupled solely via capacitors (a) or with the addition of two operational amplifiers (c). Time traces of the two input voltages (blue and orange) representing the four consecutive binary states 10, 00, 01, 11. The resulting quasi-binary output voltage (green), which is either quiescent or oscillating, is shown in (b) for circuit (a), revealing XOR behavior, and in (d) for circuit (c), exhibiting NAND behavior. The corresponding truth tables are shown next to the circuit diagrams.
To use the output of one logic gate as the input for another, the output must be compatible with the input requirements, in this case static signals of $0\,\mathrm{V}$ and $1\,\mathrm{V}$. To achieve a proper NAND gate, we must transform the oscillatory output of our NAND-pre-gate into a constant output of $1\,\mathrm{V}$ while its quiescent (0) output for the 11 input remains unaffected. This transformation is achieved by passing the output voltage of our NAND-pre-gate through the coupler shown in the circuit diagram of Fig. 6. In this coupler, the output voltage first passes through the operational amplifier OA1, which strengthens the signal and prevents drawing any current from the input gates. The amplified signal is then passed through a full-wave bridge rectifier composed of four diodes, followed by time-averaging with a low-pass filter with time constant $R_{LP}C_{LP}=0.5\,\mathrm{ms}$. Finally, the signal is amplified by OA2 to achieve the desired $1\,\mathrm{V}$ logic level.
Figure 6: Coupler for converting an oscillatory signal into constant logic levels. The oscillatory input voltage difference between $U_{\mathrm{in}}^{+}$ and $U_{\mathrm{in}}^{-}$ is first amplified by the left operational amplifier OA1. This amplified signal is then rectified by a full-wave bridge rectifier, consisting of four diodes. The rectified signal is subsequently smoothed by a low-pass filter consisting of $R_{\mathrm{LP}}$ and $C_{\mathrm{LP}}$ and finally amplified to the desired logic level by the right operational amplifier OA2. The operational amplifiers in this circuit are modeled by Eq. 3, with a maximum voltage of $U_{+}=U_{\mathrm{cc}}=1\,\mathrm{V}$.
@@
where $G$ is the gain factor (for small inputs), set to $G=15$ for $\mathrm{OA}_{1}$ and to $G=35$ for $\mathrm{OA}_{2}$. The parameter $\Delta U=U_{+}-U_{-}=1.0\,\mathrm{V}$ defines the range of the amplifier, allowing amplification up to ±$\pm 1\,\mathrm{V}$. The low-pass RC filter consists of a resistance $R_{\mathrm{LP}}=10^{-2}\,\mathrm{g_{0}}^{-1}$ and a capacitance $C_{\mathrm{LP}}=5\times 10^{-2}\,\mathrm{g_{0}s}$, while the diodes are modeled with a saturation current of $0.3\,\mathrm{g_{0}V}$ and an emission factor $\eta=1$ [47].
With the addition of this coupler circuit, the NAND-pre-gate becomes a true NAND gate.
@@ V Combining gates
The NAND gate is well known to be functionally complete, which means that any other logic gate can be constructed from it. In particular, an OR gate can be constructed using three NAND gates, as shown in Fig. 7(a). The Boolean logic for the OR gate follows $\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{1}}\cdot\overline{U_{\mathrm{in}}^{2}\cdot U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}}\cdot\overline{U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}}}+\overline{\overline{U_{\mathrm{in}}^{2}}}=U_{\mathrm{in}}^{1}+U_{\mathrm{in}}^{2}$, where $\cdot$ denotes the logic AND, $+$ the logic OR, and an overline indicates negation. In our circuits, each input signal $U_{\mathrm{in}}^{1}$ and $U_{\mathrm{in}}^{2}$ is passed through the combined gates twice to ensure independent noise sources for all our Memriki gates. If identical signals were used, the differential input to the final gate would cancel out, preventing the oscillations from self-exciting. To further aid self-excitation, we also add Gaussian noise with a standard deviation of $0.003\,\mathrm{V}$ to the inputs of the last gate.
@@
As is well known, the circuit for the OR gate can be reused to realize an AND gate; only the inputs need to be rearranged such that both input gates now receive $U_{\mathrm{in}}^{1}$ and $U_{\mathrm{in}}^{2}$, as shown in Fig. 7(b). The logic for the AND gate is as follows: $\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}\cdot\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}+\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}=U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}$. The resulting input-output relation for this AND gate is shown in Fig. 7(d), confirming that it functions correctly as an AND gate. The dynamic response to slowly varying inputs is comparable to that of the OR gate.
Figure 7: Three combined NAND gates (see text) form (a) an OR gate and (b) an AND gate. Time traces of the input voltages (blue and orange), representing four consecutive binary states, and the corresponding output voltage (green) for (c) the OR gate and (d) the AND gate, along with the corresponding truth tables. The delay in the switching of the output signal after the inputs have changed originates from the low-pass filter in the coupling circuit in  Fig. 6.
To test the repeatability and reproducibility of the gates we need to verify whether they perform as expected over many cycles of the (weakly noisy) input and output sequences. For our continuous system, with output voltages spanning the full range between zero and one volt when the input signal switches, we need to define boundaries separating high- and low-level signals during specific time intervals. Here we set the time interval as the duration between the switching of the input signals and consider an output larger than $0.66\,\mathrm{V}$ as a digital 1 and below $0.33\,\mathrm{V}$ a digital 0. To evaluate the pre-gates, we connect them to the coupling circuit as described above to ensure they function as proper gates. Using these definitions, we tested 100 cycles of all possible input combinations and obtained the following accuracies. Individual XOR gates and NAND gates operate with an accuracy of $\geq 99\%$, the combined OR gate achieves also $\geq 99\%$ accuracy, and the combined AND gate reaches an accuracy of $\geq 95\%$.
@@ VI Conclusion
We designed microfluidic iontronic circuits, primarily combining volatile memristors with Shinriki-like oscillators, that can perform non-linear XOR classification and act as functional complete NAND logic gates. We showed that these NAND gates are robust and can be combined to construct other logic gates, especially OR and AND. A key feature of these circuits is their exploitation of noise, here introduced at a level of $3\,\mathrm{mV}$ relative to a $1\,\mathrm{V}$ signal scale. This noise level enables the self-excitation of oscillators from a quiescent (0) to an oscillating (1) state, a feature that stems from the intrinsic non-linearity of the “Memriki” oscillators from which our circuits were composed. We hope our findings will inspire future experimental exploration of memristive iontronic devices. Although we have not yet assessed the energy efficiency of our circuits and logic gates in detail, we do not expect them to outperform present-day silicon-based technologies in this regard. In this sense, this work serves as a proof of concept for a potentially versatile iontronic computing platform, operating with the same aqueous electrolyte environment as mammalian brains.
@@ VII Methods
We stress that this explicit expression for $g_{\infty}(U)$ stems from an (approximate) analytic solution to the fully microscopic Poisson-Nernst-Planck-Stokes equations in the long-channel limit. The dependence of the conductance on $U$ is a direct consequence of electro-osmotic flow, characterized by the dimensionless parameter $\mathrm{Pe}(U)$. This effect relies on the presence of a nonzero surface charge $e\sigma$ (and hence a nonzero zeta potential $\psi$ on the channel walls), as becomes explicit by the analytic relation $\mathrm{Pe}(U)=\frac{-eR_{b}}{k_{B}TR_{t}w}U$, where $w=\frac{eD\eta}{k_{B}T\varepsilon\psi}\approx-9.6$ represents the ratio of ionic mobility $D/k_{B}T$ to osmotic mobility $\varepsilon\psi/(e\eta)$. Here $D=1.75\,\mathrm{nm}^{2}/\mathrm{ns}$ is the diffusion coefficient of the monovalent cations and anions in water, $k_{B}$ denotes the Boltzmann constant, $T=293.15\,\mathrm{K}$ is the absolute temperature, and $\varepsilon=80.2\epsilon_{0}\approx 7.1\times 10^{-8}\,\mathrm{F/m}$ is the dielectric constant, and $\eta=1.01\,\mathrm{mPas}$ is the shear viscosity of water. We define the steady zero-field conductance as $g_{0}=\frac{\pi R_{t}R_{b}}{L}\frac{2\rho_{b}e^{2}D}{k_{B}T}$, where $\rho_{b}$ is the ionic bulk concentration, set to $10\,\mathrm{mM}$. The asymmetry parameter is given by $\Delta g=\frac{-2w\Delta R}{R_{b}\mathrm{Du}}\approx-3.59$, where the Dukhin number is $\mathrm{Du}=\sigma/(2\rho_{b}R_{t})\approx-0.25$, with $e\sigma=-0.0015\,e/\mathrm{nm}^{2}$ the surface charge density on the channel walls. The corresponding zeta potential is $\psi=\frac{2k_{B}T}{e}\mathrm{asinh}(2\pi\lambda_{D}\lambda_{B}\sigma)\approx-10\,\mathrm{mV}$, where $\lambda_{D}=\sqrt{\varepsilon k_{B}T/(2e^{2}\rho_{b})}\approx 30\,\mathrm{nm}$ is the Debye length and $\lambda_{B}=e^{2}/(4\pi\varepsilon k_{B}T)\approx 0.71\,\mathrm{nm}$ is the Bjerrum length.